what is equivalent weight of CaSo4.1/2H2O and how to calculate it?
 

Grade:11

1 Answers

Arun
4 years ago
Molecular mass of CaSO4.1/2 H2O= Mass of one atom of calcium+ Mass of one atom of sulphur + mass of four atoms of oxygen + 9 = 40 amu + 32 amu + (4 × 16) amu + 9 amu = ( 40 + 32 + 64 + 9) amu = 145 amu.
 
for equivalent weight = 145/ change in oxidation number